# Game Theory and Multi-Agent Reinforcement Learning for Zonal Ancillary Markets

### v3

@@ V Simulation Results
We now discuss the numerical results of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary service markets, for the period between January 2024 and August 2024 [22]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 0 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other. All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory.
@@ V-A Convergence of the Multi-Agent RL Simulation
We start by analyzing the convergence and stability of the multi-agent simulation. Our settings are particularly complicated, since the market environment is dependent on all agents decisions. We tested two settings for the MARL algorithm: one with 4 learning based producers, and one with 8. In the first case, the 4 other producers always bid their marginal price. The second case creates a very dynamic environment, since at every step the price of the market is decided based on the actions of 8 independent RL agents. To assess the stability and convergence, we consider the reward of each agent throughout the training. More precisely, we compute the average reward during an episode, this being a full pass through the data. With this data we obtain Fig. 2, where we plot the episodic average reward for each agent divided by zone, considering the case with only RL based producers. Notice that the reward is not the profit, as we mention in the full description of the algorithm in the Appendix. From Fig. 2 we can see that each agent increases its reward during the initial episodes and eventually plateaus to a stable value. These results are particularly interesting, since they show how the algorithm is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
We want to stress here that, using the MARL algorithm, we compute a stationary solution for the market, without a guarantee of it being an equilibrium.
V-B MARL Solution and Classical Solutions
We first discuss the theoretical difference between the integrated solution and the Best-Response solution. Starting from the potential problem, we consider the role of the stationarity condition coming from Eq. (10), which for player $n$, bid $k$ and zone $z$ is the following:
$\begin{split}&\nabla_{x_{n,k}^{z}}L(x;\lambda,\mu,\sigma,\delta)=\Delta_{n,k}\pi_{n,k}-\lambda_{z}\Delta_{n,k}+\mu_{n,k}+\\ &+(1-\mathbb{I}_{z=z(n)})\sigma_{z(n)}\Delta_{n,k}-\mathbb{I}_{z=z(n)}\delta_{z(n)}\Delta_{n,k}.\end{split}$ (17)
Notice that every $\mu$ must be always 0, since every other term is multiplied by $\Delta$ and can be grouped. Furthermore, $\delta$ will be 0 for most realistic cases, since the core portion is usually a small fraction of the total demand, meaning that the constraint is always saturated. $\sigma$ on the other hand may be different from 0, which means that $\lambda$ must be different from 0 too, given the signs in the equation. Lastly, we highlight that $\pi$ is indexed by agent and bid, while $\lambda$ by zone, which means that, in most cases, the prices have to be equal to $\lambda$, virtually creating a pay-as-clear market. On the other hand, when using the Best Response algorithm, every agent computes its own $\lambda$, meaning that they are not forced to bid at the same price, and they can actually bid at very high prices with no repercussions. This constraint creates two very different markets, providing two different solutions to analyze.
Regarding the solution obtained from the MARL algorithm, even if there are less theoretical guarantees, we can consider it as a more realistic simulation of the market, since every producer has access only to its own information and its only aim is to increase its profit.
@@ V-B MARL Solution and Classical Solutions
We start by analyzing the average cost to run the market. This value is computed by summing the average profits of each producer over the full data available. In particular, for the MARL producers, we consider the cost for the final episode of their training, where we assume they reached stable bidding strategies. The results are reported in Fig. 3. We can immediately see how the two exact algorithms produce higher costs than the learning-based counterpart, which can be explained by the form of the constraints, as detailed above, and by the lack of actual competition. In the integrated solution, the problem is solved by an exogenous external coordinator, so by definition there will be no competition; while for the best response, each producer is free to update its decision variables uncoupled from the others. We can see the average cost for the MARL simulation with only 4 learning producers is very low, but this is also affected by the presence of static producers that always bid at marginal cost, driving the electricity price down.
@@
It is interesting to analyze how the profit is shared between producers in the different solutions. We do so by computing the Gini index [23], which, given a set of profits $p_{1}\leq p_{2}\leq\dots\leq p_{n}$, can be obtained as:
$g=\frac{1}{n}\left(n+1-2\left(\frac{\sum_{i=1}^{n}(n+1-i)p_{i}}{\sum_{i=1}^{n}p_{i}}\right)\right).$
The lower $g$ is, the more equal is the sharing of profits. We report the results of this analysis in Tab. IV. We can see that, even if the BR and potential solutions produce higher costs, they also obtain a better overall equality index, with the two MARL based simulations getting the highest index. Once again this can be explained by the fact that the multi-agent simulation is unregulated: each learning agent can bid anything it wants, which creates a situation where few producers can drive others’ profit almost to zero, using their market power. Looking at the index computed by zone, the differences between classical and learning-based algorithm stay the same.
TABLE IV: We report here the Gini index for all simulations: lower is a more equal share of the total profit between producers.
Simulation Overall Germany Austria
Best Response $0.43$ $0.16$ $0.15$
@@
We conclude with a comparison of the computational time required by each algorithm to simulate the market for the full dataset. In Tab. V we report the Evaluation and Training time: the two classical algorithms obviously have no training time, while for the MARL algorithm, the latter is what takes the longest. Although, it is important to stress that the MARL algorithm needs to be trained once and then it can be used to evaluate multiple settings without retraining, or eventually with a much shorter training routine. We also highlight that the BR algorithm’s evaluation time is longer than the training of 8 RL agents, due to the generally slow convergence.
@@ V-C Effects of Coupling Between Zones
We now analyze how the two zones influence each other in our model. To do so we simulate the market with a range of different values for the export limit $E_{z}$ in Eq. (5). We write the export limit for Germany (Austria) as a fraction of the maximal demand of Austria (Germany): $E_{G}=c_{G}\cdot\max{D_{A}}$, and we explore the ranges $c_{G}\in[0,2]$ and $c_{A}\in[0,0.2]$, so we go from completely uncoupled zone to a situation where the larger zone (Germany) has an export limit which is double the total demand of the smaller zone (Austria).
To quickly evaluate all of these values we use the MARL simulations, in both cases (4 and 8 RL agents). We report in Fig. 4 the average cost to run the market, computed in the same way as for Fig. 3, for the range of constraints values. In particular, we trained the agents for the two cases with $c_{G}=0.4,c_{A}=0.04$, and used them on all other values. When analyzing these results, the actual costs are not particularly relevant, as they depend on the demand and other external factors, while the trend of increase/decrease is what we have to interpret.
@@
We can see that in the case of $4$ RL agents, the average cost is not influenced by the export constraint: this is likely due to the fact that the presence of static agents, always bidding at marginal price, already drives the cost down, and it is not possible to decrease it further. The case with $8$ RL agents is much more interesting: we observe that the cost for Austria is rather stable, while the one for Germany is basically linearly decreasing. We can interpret this behavior considering the differences between the producers in Austria and Germany: in the smaller zone we have less producers, with lower capacities and marginal prices, while in the larger Germany the producers have higher marginal prices. Once the Austrian producers can enter the German market with their leftover capacity, they can bid the same way they would bid for their local market, which in turns, forces the German producers to bid lower, in order to be more competitive. Furthermore, since it is the market that decides how to divide each bid between zone, a producer can not create a priori a specific bid for a specific zone, each bid needs to be general enough to be competitive on every zone available. We conclude by mentioning [24], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone. Their results show the same trend we observe here, with the smaller zone obtaining similar cost, while the larger one decreases its own. Our main methodological advantage, using a MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs. Further, using pre-trained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt. This approach is of particular interest for situations where the coupling may change due to grid failures, with no warning.
@@ VI Conclusions
In this work we presented a model for multi-zonal ancillary markets based on a bilevel formulation. Each electricity producer is represented by an optimization problem with local and coupling constraints, with the objective of maximizing its profit, while the market operator objective is to minimize the total cost, so we create an adversarial formulation. We then reformulated the bilevel problem as a single level optimization, from the perspective of each producer, using the KKT reformulation of the market optimization. We interpret the ensemble of single level problems of each producer as a generalized Nash game. The game we obtain has side constraints, trilinear objective functions and nonconvex constraints, making it particularly challenging to solve. We analyze this game considering a penlaized version, and we obtain existence results for the Nash equilibrium that we relate to the initial generalized Nash game. We conclude our theoretical analysis showing that the generalized Nash game has a generalized potential structure.
We then consider different algorithms specifically tailored to our problem: as baselines, we rely on an integrated optimization reformulation relying on the potential structure of the game, and Gauss-Seidel Best Response to solve the generalized Nash game. Further, we implement a multi-agent deep reinforcement learning algorithm to simulate the market in a more realistic, distributed way.
We conclude validating the three algorithmic solutions on real data from the ancillary market between Germany and Austria. Our results showcase that computationally, the MARL simulation is the fastest to evaluate the data, but the agents have to be trained once before, while the best response approach is the slowest. The two classical algorithms also produce solutions with market costs much higher than the multi-agent counterpart, but the latter produces a more unfair split of the profits between the players. Finally, we test different forms of coupling between the zones, showing that a stronger coupling could lead to a decrease in cost for the larger zone, a result which is validated by other analysis present in the literature [24].
Appendix A KKT Reformulation Details
The full form of the Lagrangian for the market clearing problem of Eq. (8) is the following:
$\begin{split}L(x;&\lambda,\mu,\sigma,\delta)=\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}\right)\Delta_{n,k}\pi_{n,k}\\ +&\sum_{z\in\mathcal{Z}}\lambda_{z}\left(D_{z}-\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}x_{n,k}^{z}\Delta_{n,k}\right)\\ +&\sum_{n\in\mathcal{N}}\sum_{k=1}^{K_{n}}\mu_{n,k}\left(\sum_{z\in\mathcal{Z}}x_{n,k}^{z}-1\right)\\ +&\sum_{z\in\mathcal{Z}}\sigma_{z}\left(\sum_{n\in\mathcal{N}(z)}\sum_{k=1}^{K_{n}}\left(\sum_{z^{\prime}\neq z}x_{n,k}^{z^{\prime}}\right)\Delta_{n,k}-E_{z}\right)\\ +&\sum_{z\in\mathcal{Z}}\delta_{z}\left(C_{z}-\sum_{n\in\mathcal{N}(z)}x_{n,k}^{z}\Delta_{n,k}\right),\end{split}$ (18)
where we explicit the dual variables $\lambda\in\mathbb{R}_{\geq 0}^{Z},\mu\in\mathbb{R}_{\geq 0}^{B},\sigma\in\mathbb{R}_{\geq 0}^{Z},\delta\in\mathbb{R}_{\geq 0}^{Z}$.
Appendix B Penalized Objective Function
We report here the explicit form for the penalized objective function used in the penalized GNG of Eqs. (14). As described in Section III-A, we relax constraints Eqs. (3)-(6), to obtain:
$J_{n}^{p}(w)$ $=\sum_{1\leq k\leq K}\sum_{z\in\mathcal{Z}}\pi_{n,k}\Delta_{n,k}x_{n,k}^{z}$ (19a)
$+\sum_{z\in\mathcal{Z}}M_{D_{z}}\left(\sum_{n\in\mathcal{N}}\sum_{1\leq k\leq K}x_{n,k}^{z}\Delta_{n,k}-D_{z}\right)$ (19b)
$+M_{C}\left(\sum_{j\in z(n)}\sum_{0\leq k\leq K}x_{j,k}^{z(n)}\Delta_{j,k}-C_{z(n)}\right).$ (19c)
Appendix C Proof of Proposition 1
We call the two Lagrangian functions associated to the problems $\mathcal{L}_{n}^{P}$ and $\mathcal{L}_{n}^{U}$, and we consider the solution for $U$: $(w_{n}^{*},\Theta_{n}^{*})\in\mbox{SOL}\{\mbox{KKT of }U\}$, where $\Theta$ is the set of Lagrangian multipliers. This solution must respect all the constraints and be a solution of $\nabla\mathcal{L}^{U}(w^{*}_{n},\Theta^{*}_{n})=0$. We now call the part of Lagrangian multipliers that are still present in $P$ $\Theta[\Sigma_{n}]$, and we consider the set of equations coming from $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma_{n}];M_{n})=0$, where $M_{n}$ is the set of penalties for agent $n$. If we evaluate these equations in $(w_{n}^{*},\Theta_{n}^{*}[\Sigma_{n}])$, we can solve them w.r.t. the penalties $M_{n}$, obtaining $\bar{M}_{n}$. Given the assumptions, solving the system $\nabla\mathcal{L}^{P}(w_{n},\Theta_{n}[\Sigma];\bar{M}_{n})=0$ in $(w_{n},\Theta_{n}[\Sigma])$ provides $\tilde{w}_{n}$ which is a solution also for $\nabla\mathcal{L}^{U}(w_{n},\Theta_{n})=0$.
Appendix D Hessians Analysis
For the stationarity condition of Eq. (12c) we obtain the following cases:
Given that an interior point forces $\lambda_{n}^{z}=0,\delta_{n}=0$ and $\sigma_{n}=0$, and that the prices $\pi_{n,k}$ are positive by definition, both inequalities are always satisfied.
Regarding the Hessian coming from Eq. (5), which we call $\nabla^{2}G_{D_{z}}$, we can observe that deriving twice we get only constant terms, and given that all the terms in the sum are positive, we obtain a Hessian matrix with only $1$s. Then, since our decision variables are all positive by definition, the requirement $x^{\top}\nabla^{2}G_{D_{z}}x\geq 0$ is always true.
Appendix E MADDPG Implementation Details
There are three main problems related to the use of RL algorithms: defining the input state and the output action, mapping the action to satisfy the constraints and deciding the reward function.
State and Action
We choose the following forms for the input state and output action, for each producer:
  • state: $s^{t}_{z}=[p^{t-1},d^{t-1},w^{t-1},\sigma^{t-1}_{z}]$ where $p^{t-1}$ are the previous day clearing prices, $d^{t-1}$ is the day, $w^{t-1}$ is the day of the week and $\sigma_{z}^{t-1}$ is a random signal which is zone-specific.
  • action: $a^{t}_{i}=\left[\left[\pi_{n}^{t,0},\dots,\pi_{n}^{t,K_{n}}\right],\left[\Delta_{n}^{t,0},\dots,\Delta_{i}^{t,K_{n}}\right]\right]$, where $K_{n}$ is the number of bids each producer can submit.
Action Mapping
In our market model we have the following features:
  • Capacity Output: we want each capacity bid to be at least 5 MW and less than the maximum capacity, with also the sum being less or equal than the total capacity of the producer.
  • Price Output: each price bid must be larger than the marginal price $\underline{\pi}_{n}$ and smaller than the maximum bid $\bar{\pi}_{n}$
We tackle the two outputs separately, building a module for the capacity and one for the price. We use the following mapping:
  • Capacity Module: The output of the capacity module is a vector of length equal to the length of the bid sequence, which is passed through the soft-max function, representing the percentage of capacity to offer at each step. The benefit of this mapping is that is rather simple, with the drawback of forcing each producer to bid its whole capacity at each bid.
  • Price Module: The price module produces two values $\alpha,\beta\geq 0$, which are used to obtain the lowest and highest price of the bidding sequence as: $\pi_{0}=\min(\bar{\pi},\underline{\pi}+\min(\alpha,\beta))$ and $\pi_{K}=\min(\bar{\pi},\underline{\pi}+\max(\alpha,\beta))$. We then compute the other prices interpolating between $\pi_{0}$ and $\pi_{K}$.
Reward Design
The most direct approach for the reward function in the market environment would be to use the profit, but it does not encode important information: a producer may place a non-optimal bid, but if the demand is high enough, it will be accepted anyway; further the zones have different demands, so a producer in a certain zone may have a smaller profit just because the demand is lower. What we propose is a reward formed by two terms: a normalized zone profit and a penalty for bid prices that are higher than the clearing price, weighted by the hyperparameter $\beta$:
$r^{t}_{n}=\sum_{z\in\mathcal{Z}}\left(\frac{p_{n,z}^{t}}{d^{t}_{z}/|\mathcal{N}(z)|}-\beta\sum_{k=1}^{K_{n}}\mathbb{I}[\lambda_{z}^{t}>\underline{\pi}_{n}]\cdot(\pi_{n}^{t,k}-\lambda_{z}^{t})\right)$ (20)

### v4

@@ V Simulation Results
We now discuss the numerical results of the three algorithmic solutions we propose. We test these algorithms on real data from the German and Austrian ancillary services markets, for the period between January 2024 and August 2024 [noauthor_accueil_nodate]. The maximum length for the bid sequence is 5, and each submitted capacity must be $\geq 5$ MW. As shown in Tab. II, the demand in Germany is much higher and volatile than that in Austria. Furthermore, the export limit is fixed at $80$ MW for both zones and the core portion is $100$ MW for Austria and 0 MW for Germany. The symmetric and low export limit suggests that producers in Germany could have some impact in Austria, while the converse situation is less likely to arise.
@@
Finally, we test multiple values for the export constraints, in order to study the effect that the two zones have on each other. All simulations were done on a MacBook Pro with an M2 Pro chip and 16G of memory (code available at https://github.com/FrancescoMorri/MARL_ancillary_markets).
@@ V-A Convergence of the Multi-Agent RL Simulation
We start by analyzing the convergence and stability of the multi-agent simulation. Our settings are particularly complicated, since the market environment is dependent on all agents decisions. We tested two settings for MARL algorithm: one with 4 learning based producers, and one with 8. In the first case, the 4 other producers always bid their marginal price. The second case creates a very dynamic environment, since at every step the price of the market is decided based on the actions of 8 independent RL agents. To assess the stability and convergence, we consider the reward of each agent throughout the training. More precisely, we compute the average reward during an episode, which refers to the amount of steps needed to go over the full dataset [noauthor_accueil_nodate]. With this data we obtain Fig. 2, where we plot the episodic average reward for each agent divided by zone, considering the case with only RL based producers. Notice that the reward is not the profit, as we mentioned in Section IV-B. From Fig. 2 we can see that each agent increases its reward during the initial episodes and eventually plateaus to a stable value. These results are particularly interesting, since they show how the algorithm is well suited for a medium scale simulation with competitive learning agents, providing results that we can compare with the classical solutions.
V-B MARL Solution and Generalized Nash Equilibria
We first discuss the theoretical difference between the integrated solution and the Gauss-Seidel best-response solution. The two approaches may yield different equilibria: the integrated potential formulation leads to a v-GNE, whereas the best-response produces a GNE that may not be a v-GNE [^1] Indeed, the set of v-GNEs solutions of $\Gamma$ is included in the set of GNE solutions of $\Gamma$.. In case a v-GNE is reached, the market allocates the demand in Eq. (4) through a single price system, therefore leading to the determination of one price $\lambda$ for the constraint, obtaining a pay-as-clear market. On the contrary, in case there is no market to determine the price system associated with the coupling constraints, two producers $n$, $m$ might attribute different evaluations to the same constraint. This can lead to different valuations of the dual variables, in particular $\lambda_{n}\neq\lambda_{m}$, creating a virtual pay-as-bid market.
The solution obtained via MARL simulation is stationary with respect to the reward functions used by the agents. There is no guarantee that this is an exact equilibrium, but the aim is to simulate the agent interactions under varying information-sharing scenarios to analyze strategic behaviors and their impacts on market dynamics.
@@ V-B MARL Solution and Generalized Nash Equilibria
We first analyze the average cost to run the market, computed by summing the average profits of each producer over the full data available. In particular, for MARL producers, we consider the cost for the final episode of their training, where we assume they reached stable bidding strategies. The results are reported in Fig. 3. The two exact algorithms produce higher costs than the learning-based counterpart, which can be explained by the form of the constraints, as detailed above, and by the lack of actual competition. In the integrated solution, the problem is solved by an exogenous external coordinator, so by definition there will be no competition; while for the best response, each producer is free to update its decision variables uncoupled from the others. We can see the average cost for MARL simulation with only 4 learning producers is very low, but this is also affected by the presence of static producers that always bid at marginal cost, driving the electricity price down.
@@
It is interesting to analyze how the profit is shared between producers in the different solutions. To that purpose, we rely on the Gini index $\mathscr{G}$ [gini_measure_1936]. Given a finite sequence of profits ranked in increasing order $J^{(1)}\leq J^{(2)}\leq\dots\leq J^{(n)}$, where $J^{(l)}$ denotes the $l$-th highest profit, $\mathscr{G}$ can be obtained as:
$\mathscr{G}:=\frac{1}{n}\left(n+1-2\left(\frac{\sum_{i=1}^{n}(n+1-i)J^{(i)}}{\sum_{i=1}^{n}J^{(i)}}\right)\right).$
The lower $\mathscr{G}$ is, the more equal is the sharing of profits. We report the results of this analysis in Tab. IV. We can see that, even if the BR and potential solutions produce higher costs, they also obtain a better overall equality index, with the two MARL based simulations getting the highest index. Once again this can be explained by the fact that the multi-agent simulation is unregulated: each learning agent can bid anything it wants, which creates a situation where few producers can drive others’ profit almost to zero, using their market power. Looking at the index computed by zone, the differences between classical and learning-based algorithm stay the same.
TABLE IV: Gini index for all simulations.
Simulation Overall $\mathscr{G}$ Germany $\mathscr{G}$ Austria $\mathscr{G}$
Best-Response $0.43$ $0.16$ $0.15$
@@
We conclude with a comparison of the computational time required by each algorithm to simulate the market for the full dataset. In Tab. V we report the evaluation and training times: the two classical algorithms obviously have no training time, while this is the most time-consuming part for MARL. Further, MARL achieves faster evaluation time than the exact methods. The BR algorithm’s evaluation time is longer than the training of 8 RL agents, due to the sequential calls of the optimization solver.
@@ V-C Effects of Coupling Between Zones
We now analyze how the two zones influence each other in our model. To do so we simulate the market with a range of different values for the export limit $E_{z}$ in Eq. (6). We write the export limit for Germany (Austria) as a fraction of the maximal demand of Austria (Germany): $E_{G}=c_{G}\cdot\max{D_{A}}$, and we explore the ranges $c_{G}\in[0,2]$ and $c_{A}\in[0,0.2]$, thus we go from uncoupled zones to a scenario where the larger zone (Germany) has an export limit which is double the total demand of the smaller zone (Austria).
@@
To quickly evaluate all of these values we use the MARL simulations, in both cases (4 and 8 RL agents). We report in Fig. 4 the average cost to run the market, computed in the same way as for Fig. 3, for the range of constraints values. In particular, we trained the agents with $c_{G}=0.4,c_{A}=0.04$, and used them on all other values. We can see that in the case of $4$ RL agents, the average cost is not influenced by the export constraint: this is likely due to the fact that the presence of static agents, always bidding at marginal price, already drives the cost down, and it is not possible to decrease it further. The case with $8$ RL agents is much more interesting: we observe that the cost for Austria is rather stable, while the one for Germany is basically linearly decreasing. We can interpret this behavior considering the differences between the producers in Austria and Germany: in the smaller zone we have less producers, with lower capacities and marginal prices, while in the larger Germany the producers have higher marginal prices. Once the Austrian producers can enter the German market with their leftover capacity, they can bid the same way they would bid for their local market, which in turns, forces the German producers to bid lower, in order to be more competitive. Furthermore, since it is the market that decides how to divide each bid between zone, a producer can not create a priori a specific bid for a specific zone, each bid needs to be general enough to be competitive on every zone available. We conclude by mentioning [di_cosmo_welfare_2020], where the authors simulated a more economically sophisticated model of the European electricity market, to test what the addition of another smaller zone would do to the costs in the French zone. Their results show the same trend as ours, with the smaller zone obtaining similar cost, while the larger one decreases its own. The main advantage, using MARL based simulation, is that we can quickly test a range of different settings, not just a single scenario, providing a less detailed but broader view of the effect of different market designs. Further, using pretrained agents, we are testing how a bidding strategy developed for a certain market design translates to a different one, when there is no time to adapt. This is relevant for scenarios where the coupling may change due to grid failures, with no warning.
@@ VI Conclusions
In this work, we rely on an adversarial Stackelberg game model of the multi-zonal ancillary markets, to study the impact of zonal coupling, that we recast as a nonconvex generalized Nash game with side constraints involving an associated price clearance to be satisfied by the equilibria. Further, relying on the potential structure of the generalized Nash game, we show the existence of a Nash equilibrium. Equilibrium can be reached making use of the game structure and relying on reformulations. Specifically, taking advantage of the potential structure of the game, we compare two exact approaches: Gauss-Seidel best-response and an integrated optimization reformulation. Another paradigm relies on the implementation of a multi-agent reinforcement learning (MARL) framework to simulate decentralized agent interactions, allowing us to study strategic behaviors and their impact on market dynamics. We validate the three algorithmic approaches using real data from the ancillary market between Germany and Austria. Our results show that, computationally, the MARL simulation (counting training and evaluation time) is more efficient than the best-response method, while remaining a decentralized approach. The two exact algorithms produce solutions with significantly higher total market costs compared to MARL; although the latter results in a more unfair allocation of profits among players. Numerically, we analyze different levels of coupling between zones, highlighting that tighter coupling may reduce costs in the larger zone; a result which is consistent with findings reported in the literature, e.g., in [di_cosmo_welfare_2020].